# Smooth band termination in 115I

### v2

@@ Abstract
journal: Physics Letters B
1 Introduction
The band structures of deformed nuclei with a reduced number of protons and neutrons above the $Z,N=50$ magic gaps offer optimum test benches for theoretical models conceived to describe the interplay between collective and single-particle motion. One of these models is cranked Nilsson-Strutinsky (CNS) [1], developed in particular for the interpretation of high-spin states. As pairing is neglected in this model, it has generally not been applied to low-spin states. However, the spin dependence of the difference between observed and calculated energies should mirror the pairing energy. Indeed, if it is accepted that the difference between experiment and calculations can be understood as an average pairing energy, it appears that the model can be used for all spin values, see e.g. [2]. An exception may be the region of the first band-crossing in prolate nuclei where the Fermi level is located at the bottom of a high-$j$ shell. However, the general features of such band-crossings are well understood from simple models, see e.g. [3] or [4].
The present work is devoted to 115I, a nucleus studied in the past using $\gamma$-ray spectrometers with low and moderate efficiency [5, 6, 7, 8], but also very recently in a high-statistics experiment with the JUROGAM array [9], which allowed to extend the previously known bands and to identify several bands to spin values above II=30 [10]. The use of clover detectors in the JUROGAM experiment was beneficial, allowing the firm assignment of spins and parities to all bands from angular correlations and polarization measurements. The spins and parities of several levels reported previously have been changed, and the level scheme has been largely revised. The experimental information on the $\gamma$-ray transitions observed in 115I will be reported in a forthcoming paper [10]. In the present work we only show the bands relative to a rotating liquid drop energy in Fig. 1, and give the band-head spins and parities, the band-head energies, and the transition energies of the bands in Table 1.
Table 1: Experimental information for the bands of 115I discussed in the present work: band labels, band-head spins and parities, band-head energies (in keV), parities of the bands, and transition energies (in keV) of the in-band transitions. The new transitions, to be published in a forthcoming paper [10], are indicated wtih bold characters. For band 8, the two cascades of $E2$ transitions interconnected by dipole transitions are labeled 8e and 8o, corresponding to even and odd signature, respectively. The highest transitions of band 7 are from Ref. [7]. The energies of these transitions have not been observed in Ref. [10], therefore the energies are in parentheses.
@@ 1 Introduction
 410 1032  505 617 $\bf\leavevmode\nobreak\ 654$  736 $\bf 1028$  633  662
 516 1053  488 705 $\bf\leavevmode\nobreak\ 762$  845 $\bf 1090$  700  732
 622 $\bf 1088$  646 781 $\bf\leavevmode\nobreak\ 710$  930 $\bf 1146$  758  778
@@
 796  820 665  288  593 $\bf 1298$  797 $\bf\leavevmode\nobreak\ 823$
 849  924 818  678  686 $\bf 1382$ $\bf\leavevmode\nobreak\ 862$ $\bf\leavevmode\nobreak\ 905$
 893 $\bf\leavevmode\nobreak\ 924$  866  786 $\bf 1467$ $\bf\leavevmode\nobreak\ 938$ $\bf\leavevmode\nobreak\ 977$
1015 $\bf\leavevmode\nobreak\ 736$ $\bf 1091$  880 $\bf 1575$ $\bf 1010$ $\bf 1035$
1055 $\bf\leavevmode\nobreak\ 993$ $\bf 1285$  936 $\bf 1705$ $\bf 1071$ $\bf 1092$
@@
2 General considerations
The band structure of 115I analyzed in the present work is composed of two negative-parity bands (bands 1 and 2) and six positive-parity bands (bands 3-8). Band 1 with negative parity and four positive-parity bands start at low spin and are interpreted as built on one-quasiparticle configurations. Band 1 built on the $11/2^{-}$ state is clearly based on the $\nu h_{11/2}$ configuration. Band 3, previously reported with negative-parity [8], has in fact positive parity, as established from the polarization asymmetry of the 936-keV transition to the $27/2^{-}$ state of band 1, which indicates an $E1$ character [10]. It decays via five transitions to states of the negative-parity band 1, like in the isotone 117Cs [11], and to the positive-parity bands 4 and 6. Numerous interconnecting transitions exist between the four bands 3-6, strongly supporting similar configurations. Band 8 consisting of two nearly degenerate cascades of $E2$ transitions interconnected by $M1/E2$ dipole transitions is clearly based on a configuration with one hole in the $\pi g_{9/2}$ orbital, which give rise to strongly coupled bands, like in several nuclei of the AA=115 mass region, see e. g. [11, 12, 13, 14, 15, 16].
Once the configurations are assigned to the low-spin parts of the bands, one would like to understand their evolution at high spin. One can get a first idea by realizing the resemblance of the observed bands with those recently reported in the 117Cs isotone [11], in which the low-spin bands 1, 2, 3, 4 correspond one-to-one to the bands 1, 3, 4, 8 of 115I. This is not surprising, because the two nuclei are isotones and the number of protons is different by two.
@@ 2 General considerations
Figure 1: (Color online) (a) $E$-$E_{rld}$ values for the bands of 115I. (b) Spin II versus $\gamma$-ray energy $E_{\gamma}$. Note that bands 4, 5 and 6 show a clear backbend at $E_{\gamma}\approx 0.7$ MeV, while only a minor upbend is seen for band 1.
@@
For II>>20, the slopes of bands 3 and 6 change drastically, which makes them rapidly yrast. A special attention should be given to band 6, which becomes rapidly non yrast at the two highest states, and behaves like a smooth terminating band [17, 18].
@@
2.1 CNS calculations and configuration assignments
In the CNS formalism [1, 19], the nucleus rotates about one of its principal axes and pairing is neglected. The deformation is optimized in $(\varepsilon_{2},\gamma,\varepsilon_{4})$ for each configuration at each spin value. The configurations are labelled by the number of particles in low-$j$ and high-$j$ orbitals, in the different $\cal N$}-shells. For 115I, they can be defined relative to a 100Sn core as,
@@ 2.1 CNS calculations and configuration assignments
Figure 2: (Color online) The calculated low-energy CNS configurations in the valence space of 110I: (a) positive and (b) negative parity. States with positive (negative) signature are drawn with filled (open) symbols.
@@
Figure 3: (Color online) Deformation trajectories for the positive-parity valence space configurations in Fig. 1(a).
@@
Figure 4: (Color online) (a) Single-proton and (b) single-neutron Routhians at a prolate shape typical for 115I. The orbitals are traced back to spherical shape on the left. Line convention for $(\pi,\alpha)$ orbitals: continuous for $(+,+1/2)$, dotted for (+,1/2)(+,-1/2), dashed for (,+1/2)(-,+1/2), doted-dashed for (,1/2)(-,-1/2). The occupation of the orbitals in the terminating $21_{-};9_{+}3_{-}]$ ($\pi(dg)^{2}(h_{11/2})^{1}\nu(dg)^{9}(h_{11/2})^{3}$) configuration is illustrated by filled black circles.
2.1.1 Bands 1 and 2
@@
Figure 6: (Color online) Same as Fig. 5 but comparing of the experimental bands 3 and 6 with CNS configurations.
@@
Figure 8: (Color online) Same as Fig. 5 but comparing of the experimental bands 7 and 8 with CNS configurations.
2.1.2 Bands 3-6
@@
Here we will concentrate on bands 3 and 6 which are observed to highest spin values. These bands are assigned to $(dg)^{3}$ proton configurations combined with neutron configurations with an even number of neutrons in the positive- and negative-parity orbitals, i.e. $(dg)^{8}(h_{11/2})^{4}$ or $(dg)^{10}(h_{11/2})^{2}$ according to the discussion above. The energy curves of these configurations are drawn in the middle panel of Fig. 6 where we note a crossing at $I\approx$15, where the $(h_{11/2})^{4}$ bands which are lower in energy at low spin, are crossed by the $(h_{11/2})^{2}$ bands which are lower at high spin. This observation suggests that the sharp crossings at II=10 are created by a change in configuration from a $(h_{11/2})^{4}$ configuration to a $(h_{11/2})^{2}$ configuration. Our interpretation is supported from the analysis of bands 1 and 2, showing that the relative energies of the $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ configurations are more or less reproduced in our calculations. Furthermore, both in calculations, in general, and in experiment for bands 1 and 2, the $\nu(h_{11/2})^{2}$ configuration comes relatively lower in energy with increasing spin. Thus, the combination of these two neutron configurations with the two signatures of the proton in the $(dg)$ 3/2 [422] orbital will lead to two pairs of bands which cross sharply somewhere around II=10-15. This is in general agreement with experiment, while a detailed understanding of their configurations could not really be expected in the CNS formalism. This interpretation is in line with general expectations that in high-$j$ shells, angular momentum is built more efficiently with fewer particles in the shell. This is the case for the proton $h_{11/2}$ shell in the ZZ=66-70 nuclei where with increasing spin, the number of $h_{11/2}$ protons for the yrast states decreases [21], $(h_{11/2})^{8}\rightarrow(h_{11/2})^{6}\rightarrow(h_{11/2})^{4}$ or $(h_{11/2})^{7}\rightarrow(h_{11/2})^{5}$, see also Fig. 12.11 in Ref. [4].
Crossings in high-$j$ shells are generally understood as so called ’paired crossings’, i.e. a re-coupling of the spin vectors within the shell, see e. g. [3, 4, 22]. This is well understood in the bottom of the shell. However, as shown in Ref. [18], higher up in the shell when the pairing is not too strong, such crossings are naturally understood as a decrease of the number of particles in the high-$j$ shell. This interpretation challenges the more standard interpretation of crossings like that in bands 3-6 in 115I. However, it offers an explanation to the fact that only a very smooth crossing is observed in band 1, while a much sharper crossing is seen in bands 3-6.
@@
where the spin values in the different groups are specified in the subscripts. Note that this is the maximum spin of signature $\alpha$=1/2-1/2 in the [21;93] configurations. To our knowledge, while many smooth terminating bands with holes in the proton core have been observed in this space [17, 18], this is the first time that such a band within the valence space has been observed up to its maximum spin value, $I_{max}$.
2.1.3 Bands 7 and 8
@@
Figure 9: (Color online) (a) Low-energy configurations with two proton $g_{9/2}$ holes. (b) Low-energy configurations with one proton $g_{9/2}$ hole and an even number of particles in the other groups.
@@
The bands 7 and 8 with two holes and one hole in the $g_{9/2}$ proton shell, respectively, are compared with their calculated counterpart in Fig. 8. The difference curves come out as expected with a smooth decrease with increasing spin. Indeed, if the $g_{9/2}$ orbital was lifted a few hundred keV, the difference curve would more or less coincide with the one which is used as an average for all bands of 115I, see Fig. 8(c). Note especially that while the $E$-$E_{rld}$ curves for band 8 in Fig. 6(a) and the configurations assigned to it in Fig. 6(b) do not show much resemblance, the difference curve in Fig. 6(c) comes out as expected considering that pairing is neglected in the calculations. We have also included band 1 in 113I [14]. It is assigned to a configuration with two less $(dg)$ neutrons than the configuration of band 7 in 115I and leads to a very similar difference curve. This comparison illustrates the possibility to compare bands in different nuclei in the present CNS formalism [19] and gives some extra confidence that our assignments for bands 7 and 8 are correct.
3 Summary
@@
Also the positive-parity bands with all three protons in the $(dg)$ orbitals combine with the $(dg)^{8}(h_{11/2})^{4}$ and $(dg)^{10}(h_{11/2})^{2}$ neutron configurations. In this case, the details are more difficult to understand but it is suggested that the sharp crossings in the bands at $I\approx$1010 can be understood as a crossing between $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ and configurations. At higher spin values above $I\approx$2020 for these positive-parity bands, the $(dg)^{3}$ proton configuration is not competitive, i.e. the $(dg)^{2}(h_{11/2})^{1}$ configuration takes over, combined with the $\nu(h_{11/2})^{3}$ configuration. In this case, the two signatures lead to one smooth terminating band and another band which becomes non-collective before the $I_{max}$-value is reached. This smooth terminating band is observed up to the maximum spin state in the configuration, i.e. to the $I_{max}$ state. While many such bands with holes in the $g_{9/2}$ orbitals have been observed for nuclei with mass numbers just below 115I [17, 18], this is to our knowledge the first time that such a smooth terminating band has been observed to termination within the present valence space.

### v3

@@ Abstract
I Introduction
The band structures of deformed nuclei with a reduced number of protons and neutrons above the $Z,N=50$ magic gaps offer optimum test benches for theoretical models conceived to describe the interplay between collective and single-particle motion. One of these models is cranked Nilsson-Strutinsky (CNS) Bengtsson and Ragnarsson (1985a), developed in particular for the interpretation of high-spin states. As pairing is neglected in this model, it has generally not been applied to low-spin states. However, the spin dependence of the difference between observed and calculated energies should mirror the pairing energy. Indeed, if it is accepted that the difference between experiment and calculations can be understood as an average pairing energy, it appears that the model can be used for all spin values, see e.g. Ragnarsson et al. (2024). An exception may be the region of the first band-crossing in prolate nuclei where the Fermi level is located at the bottom of a high-$j$ shell. However, the general features of such band-crossings are well understood from simple models, see e.g. Stephens et al. (1973) or Nilsson and Ragnarsson .
The present work is devoted to 115I, a nucleus studied in the past using $\gamma$-ray spectrometers with low and moderate efficiency Piel et al. (1985); Paul et al. (1992); Paul et al. (1994); Moon et al. (2003), but also very recently in a high-statistics experiment with the JUROGAM array Pakarinen et al. (2020), which allowed to extend the previously known bands and to identify several bands to spin values above II=30 Jodidar and Petrache . The use of clover detectors in the JUROGAM experiment was beneficial, allowing the firm assignment of spins and parities to all bands from angular correlations and polarization measurements. The spins and parities of several levels reported previously have been changed, and the level scheme has been largely revised. The experimental information on the $\gamma$-ray transitions observed in 115I will be reported in a forthcoming paper Jodidar and Petrache . In the present work we only show the bands relative to a rotating liquid drop energy in Fig. 1, and give the band-head spins and parities, the band-head energies, and the transition energies of the bands in Table 1.
Table 1: Experimental information for the bands of 115I discussed in the present work: band labels, band-head spins and parities, band-head energies (in keV), parities of the bands, and transition energies (in keV) of the in-band transitions. The new transitions, to be published in a forthcoming paper Jodidar and Petrache , are indicated wtih bold characters. For band 8, the two cascades of $E2$ transitions interconnected by dipole transitions are labeled 8e and 8o, corresponding to even and odd signature, respectively. The highest transitions of band 7 are from Ref. Paul et al. (1994). The energies of these transitions have not been observed in Ref. Jodidar and Petrache , therefore the energies are in parentheses.
@@ I Introduction
410 1032 505 617 $\bf~654$ 736 $\bf 1028$ 633 662
516 1053 488 705 $\bf~762$ 845 $\bf 1090$ 700 732
622 $\bf 1088$ 646 781 $\bf~710$ 930 $\bf 1146$ 758 778
@@
796 820 665 288 593 $\bf 1298$ 797 $\bf~823$
849 924 818 678 686 $\bf 1382$ $\bf~862$ $\bf~905$
893 $\bf~924$ 866 786 $\bf 1467$ $\bf~938$ $\bf~977$
1015 $\bf~736$ $\bf 1091$ 880 $\bf 1575$ $\bf 1010$ $\bf 1035$
1055 $\bf~993$ $\bf 1285$ 936 $\bf 1705$ $\bf 1071$ $\bf 1092$
@@
II General considerations
The band structure of 115I analyzed in the present work is composed of two negative-parity bands (bands 1 and 2) and six positive-parity bands (bands 3-8). Band 1 with negative parity and four positive-parity bands start at low spin and are interpreted as built on one-quasiparticle configurations. Band 1 built on the $11/2^{-}$ state is clearly based on the $\nu h_{11/2}$ configuration. Band 3, previously reported with negative-parity Moon et al. (2003), has in fact positive parity, as established from the polarization asymmetry of the 936-keV transition to the $27/2^{-}$ state of band 1, which indicates an $E1$ character Jodidar and Petrache . It decays via five transitions to states of the negative-parity band 1, like in the isotone 117Cs Jodidar et al. (2024), and to the positive-parity bands 4 and 6. Numerous interconnecting transitions exist between the four bands 3-6, strongly supporting similar configurations. Band 8 consisting of two nearly degenerate cascades of $E2$ transitions interconnected by $M1/E2$ dipole transitions is clearly based on a configuration with one hole in the $\pi g_{9/2}$ orbital, which give rise to strongly coupled bands, like in several nuclei of the AA=115 mass region, see e. g. Jodidar et al. (2024); Evans et al. (2006); Warring et al. (1993); Starosta et al. (2001); Zheng et al. (2021); Jodidar and Petrache (2025).
Once the configurations are assigned to the low-spin parts of the bands, one would like to understand their evolution at high spin. One can get a first idea by realizing the resemblance of the observed bands with those recently reported in the 117Cs isotone Jodidar et al. (2024), in which the low-spin bands 1, 2, 3, 4 correspond one-to-one to the bands 1, 3, 4, 8 of 115I. This is not surprising, because the two nuclei are isotones and the number of protons is different by two.
@@ II General considerations
Figure 1: (a) $E$-$E_{rld}$ values for the bands of 115I. (b) Spin II versus $\gamma$-ray energy $E_{\gamma}$. Note that bands 4, 5 and 6 show a clear backbend at $E_{\gamma}\approx 0.7$ MeV, while only a minor upbend is seen for band 1.
@@
For II>>20, the slopes of bands 3 and 6 change drastically, which makes them rapidly yrast. A special attention should be given to band 6, which becomes rapidly non yrast at the two highest states, and behaves like a smooth terminating band Ragnarsson et al. (1995); Afanasjev et al. (1999).
@@
II.1 CNS calculations and configuration assignments
In the CNS formalism Bengtsson and Ragnarsson (1985a); Carlsson and Ragnarsson (2006), the nucleus rotates about one of its principal axes and pairing is neglected. The deformation is optimized in $(\varepsilon_{2},\gamma,\varepsilon_{4})$ for each configuration at each spin value. The configurations are labelled by the number of particles in low-$j$ and high-$j$ orbitals, in the different $\cal N$}-shells. For 115I, they can be defined relative to a 100Sn core as,
@@ II.1 CNS calculations and configuration assignments
Figure 2: The calculated low-energy CNS configurations in the valence space of 110I: (a) positive and (b) negative parity. States with positive (negative) signature are drawn with filled (open) symbols.
@@
Figure 3: Deformation trajectories for the positive-parity valence space configurations in Fig. 1(a).
@@
Figure 4: (a) Single-proton and (b) single-neutron Routhians at a prolate shape typical for 115I. The orbitals are traced back to spherical shape on the left. Line convention for $(\pi,\alpha)$ orbitals: continuous for $(+,+1/2)$, dotted for (+,1/2)(+,-1/2), dashed for (,+1/2)(-,+1/2), doted-dashed for (,1/2)(-,-1/2). The occupation of the orbitals in the terminating $21_{-};9_{+}3_{-}]$ ($\pi(dg)^{2}(h_{11/2})^{1}\nu(dg)^{9}(h_{11/2})^{3}$) configuration is illustrated by filled black circles.
II.1.1 Bands 1 and 2
@@
Figure 6: Same as Fig. 5 but comparing of the experimental bands 3 and 6 with CNS configurations.
@@
Figure 8: Same as Fig. 5 but comparing of the experimental bands 7 and 8 with CNS configurations.
II.1.2 Bands 3-6
@@
Here we will concentrate on bands 3 and 6 which are observed to highest spin values. These bands are assigned to $(dg)^{3}$ proton configurations combined with neutron configurations with an even number of neutrons in the positive- and negative-parity orbitals, i.e. $(dg)^{8}(h_{11/2})^{4}$ or $(dg)^{10}(h_{11/2})^{2}$ according to the discussion above. The energy curves of these configurations are drawn in the middle panel of Fig. 6 where we note a crossing at $I\approx$15, where the $(h_{11/2})^{4}$ bands which are lower in energy at low spin, are crossed by the $(h_{11/2})^{2}$ bands which are lower at high spin. This observation suggests that the sharp crossings at II=10 are created by a change in configuration from a $(h_{11/2})^{4}$ configuration to a $(h_{11/2})^{2}$ configuration. Our interpretation is supported from the analysis of bands 1 and 2, showing that the relative energies of the $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ configurations are more or less reproduced in our calculations. Furthermore, both in calculations, in general, and in experiment for bands 1 and 2, the $\nu(h_{11/2})^{2}$ configuration comes relatively lower in energy with increasing spin. Thus, the combination of these two neutron configurations with the two signatures of the proton in the $(dg)$ 3/2 [422] orbital will lead to two pairs of bands which cross sharply somewhere around II=10-15. This is in general agreement with experiment, while a detailed understanding of their configurations could not really be expected in the CNS formalism. This interpretation is in line with general expectations that in high-$j$ shells, angular momentum is built more efficiently with fewer particles in the shell. This is the case for the proton $h_{11/2}$ shell in the ZZ=66-70 nuclei where with increasing spin, the number of $h_{11/2}$ protons for the yrast states decreases Bengtsson and Ragnarsson (1985b), $(h_{11/2})^{8}\rightarrow(h_{11/2})^{6}\rightarrow(h_{11/2})^{4}$ or $(h_{11/2})^{7}\rightarrow(h_{11/2})^{5}$, see also Fig. 12.11 in Ref. Nilsson and Ragnarsson .
Crossings in high-$j$ shells are generally understood as so called ’paired crossings’, i.e. a re-coupling of the spin vectors within the shell, see e. g. Stephens et al. (1973); Nilsson and Ragnarsson ; Frauendorf (1981). This is well understood in the bottom of the shell. However, as shown in Ref. Afanasjev et al. (1999), higher up in the shell when the pairing is not too strong, such crossings are naturally understood as a decrease of the number of particles in the high-$j$ shell. This interpretation challenges the more standard interpretation of crossings like that in bands 3-6 in 115I. However, it offers an explanation to the fact that only a very smooth crossing is observed in band 1, while a much sharper crossing is seen in bands 3-6.
@@
where the spin values in the different groups are specified in the subscripts. Note that this is the maximum spin of signature $\alpha$=1/2-1/2 in the [21;93] configurations. To our knowledge, while many smooth terminating bands with holes in the proton core have been observed in this space Ragnarsson et al. (1995); Afanasjev et al. (1999), this is the first time that such a band within the valence space has been observed up to its maximum spin value, $I_{max}$.
II.1.3 Bands 7 and 8
@@
Figure 9: (a) Low-energy configurations with two proton $g_{9/2}$ holes. (b) Low-energy configurations with one proton $g_{9/2}$ hole and an even number of particles in the other groups.
@@
The bands 7 and 8 with two holes and one hole in the $g_{9/2}$ proton shell, respectively, are compared with their calculated counterpart in Fig. 8. The difference curves come out as expected with a smooth decrease with increasing spin. Indeed, if the $g_{9/2}$ orbital was lifted a few hundred keV, the difference curve would more or less coincide with the one which is used as an average for all bands of 115I, see Fig. 8(c). Note especially that while the $E$-$E_{rld}$ curves for band 8 in Fig. 6(a) and the configurations assigned to it in Fig. 6(b) do not show much resemblance, the difference curve in Fig. 6(c) comes out as expected considering that pairing is neglected in the calculations. We have also included band 1 in 113I Starosta et al. (2001). It is assigned to a configuration with two less $(dg)$ neutrons than the configuration of band 7 in 115I and leads to a very similar difference curve. This comparison illustrates the possibility to compare bands in different nuclei in the present CNS formalism Carlsson and Ragnarsson (2006) and gives some extra confidence that our assignments for bands 7 and 8 are correct.
III Summary
@@
Also the positive-parity bands with all three protons in the $(dg)$ orbitals combine with the $(dg)^{8}(h_{11/2})^{4}$ and $(dg)^{10}(h_{11/2})^{2}$ neutron configurations. In this case, the details are more difficult to understand but it is suggested that the sharp crossings in the bands at $I\approx$1010 can be understood as a crossing between $\nu(h_{11/2})^{4}$ and $\nu(h_{11/2})^{2}$ and configurations. At higher spin values above $I\approx$2020 for these positive-parity bands, the $(dg)^{3}$ proton configuration is not competitive, i.e. the $(dg)^{2}(h_{11/2})^{1}$ configuration takes over, combined with the $\nu(h_{11/2})^{3}$ configuration. In this case, the two signatures lead to one smooth terminating band and another band which becomes non-collective before the $I_{max}$-value is reached. This smooth terminating band is observed up to the maximum spin state in the configuration, i.e. to the $I_{max}$ state. While many such bands with holes in the $g_{9/2}$ orbitals have been observed for nuclei with mass numbers just below 115I Ragnarsson et al. (1995); Afanasjev et al. (1999), this is to our knowledge the first time that such a smooth terminating band has been observed to termination within the present valence space.